# CUHK-EE Systems for the vTAD Challenge at NCMMSC 2025

### v1

@@ Abstract
This paper presents the Voice Timbre Attribute Detection (vTAD) systems developed by the Digital Signal Processing & Speech Technology Laboratory (DSP&STL) of the Department of Electronic Engineering (EE) at The Chinese University of Hong Kong (CUHK) for the 20th National Conference on Human-Computer Speech Communication (NCMMSC 2025) vTAD Challenge. The proposed systems leverage WavLM-Large embeddings with attentive statistical pooling to extract robust speaker representations, followed by two variants of Diff-Net, i.e., Feed-Forward Neural Network (FFN) and Squeeze-and-Excitation-enhanced Residual FFN (SE-ResFFN), to compare timbre attribute intensities between utterance pairs. Experimental results demonstrate that the WavLM-Large+FFN system generalises better to unseen speakers, achieving 77.96% accuracy and 21.79% EER, while the WavLM-Large+SE-ResFFN model excels in the ‘Seen’ setting with 94.42% accuracy and 5.49% EER. These findings highlight a trade-off between model complexity and generalisation, and underscore the importance of architectural choices in fine-grained speaker modelling. Our analysis also reveals the impact of speaker identity, annotation subjectivity, and data imbalance on system performance, pointing to future directions for improving robustness and fairness in timbre attribute detection.
@@ 1 Introduction
The design of CUHK-EE systems entering the vTAD 2025 leverages a large-scale speech self-supervised learning (SSL) representation model, namely the WavLM-Large [3], to extract robust voice features from speech signals. The voice features are refined using attentive statistical pooling (ASTP) [4] before being passed into a comparison network, referred to as the Diff-Net. Two architectural variants of Diff-Net are investigated: standard Feed-Forward Neural Network (FFN) and Deep Residual FFN enhanced with Squeeze-and-Excitation blocks (SE-ResFFN) [5, 6, 7].
@@ 2.2 System Overview
The Diff-Net module outputs a vector of dimension $N$. Each element of the output vector represents a specific timbre attribute. A sigmoid function is applied to the output of the Diff-Net to produce the prediction vector $\hat{\rm y}$. The $n$-th element of $\hat{\rm y}$ ($n=1,2,...,34$) gives the likelihood that ${\mathcal{O}}_{\rm B}$ has a stronger intensity than ${\mathcal{O}}_{\rm A}$ in the $n$-th timbre descriptor.
@@ 2.4 Diff-Net
Figure 3 depicts the architectures of the two variants of Diff-Net investigated in our study. The first variant, illustrated as in Figure 3 (a), is an enhanced version of the traditional FFN [1, 2]. It consists of four sequential FFN blocks with hidden dimensions of [512,256,128,64][512,256,128,64][ 512 , 256 , 128 , 64 ].. Each block comprises a fully connected (FC) layer, followed by batch normalisation (BN), a ReLU activation function, and a dropout layer with a dropout rate of 0.3. After the final FFN block, an additional FC layer is applied to produce predictions for all 34 timbre attributes.
The second variant, SE-ResFFN, is inspired by the squeeze-and-excitation ResNet (SE-ResNet) architecture, which has shown strong performance in SV tasks [5, 7]. As illustrated in Figure 3 (b), (c), and (d), this model incorporates four SE-ResNet blocks with hidden dimensions of [1024,1024,512,256][1024,1024,512,256][ 1024 , 1024 , 512 , 256 ]. Following these blocks, a BN layer is applied, and the resulting features are passed through two FC layers with hidden dimensions [192,64][192,64][ 192 , 64 ], interleaved with a ReLU activation. A final FC layer then outputs the predictions for the 34 timbre attributes.
@@ 3.1 Datasets
The VCTK-RVA dataset [12] is used for training the downstream Diff-Net modules (i.e., FFN and SE-ResFFN) and performance evaluation while the core component of the speaker encoder, the pre-trained WavLM-Large, is frozen in our training procedures. VCTK-RVA provides the annotation of 17 timbre descriptors for male and female speakers, resulting in a total of 34 distinct timbre attributes. There are over 6,000 annotated speaker pairs in the dataset. Each pair is labelled with one to three timbre descriptors. The descriptors characterise different aspects of perceived voice quality, such as bright, thin, coarse, magnetic, shrill, and husky. The annotations are derived from human judgment, reflect subjective interpretations of vocal timbre. The subjectivity introduces variability in the data, which poses challenges for model consistency and generalisation to unseen speakers and utterances.
@@ 3.2 Training Settings
Model training was conducted over 10 epochs with a batch size of 16, using the AdamW optimiser with a learning rate of $1e-4$ and a weight decay of 0.01. A cosine annealing scheduler was employed to adjust the learning rate dynamically throughout training. The loss function used was binary cross-entropy with sample-wise reduction, and model validation was performed after each epoch. To ensure reproducibility, all experiments were initialised with a fixed random seed (42).
@@
Table 1: The results on the test set.
@@
Seen WavLM-Large +SE-ResFFN 94.42 5.49
@@ 4 Results and Analysis
To further probe model robustness, we conducted small-scale experiments using speaker-disjoint training splits and observed considerable variation in results depending on the speaker composition. This, along with the inherent subjectivity in manual timbre labelling, suggests that model performance is sensitive to both speaker identity and annotation consistency. Additionally, the dataset [12] exhibits imbalance in descriptor frequency and gender representation–some attributes are under-represented, and female speakers dominate the data, yet male speakers often yield better results. These factors collectively highlight the need for more balanced data and refined annotation practices to improve generalisation and fairness.
Detailed results for the ‘Unseen’ and ‘Seen’ test sets, broken down by gender and descriptor, are listed in Table 2.
Table 2: Evaluation results of our proposed systems on the ‘Unseen’ and ‘Seen’ test sets. The row ‘Avg’ is obtained by averaging the results across all the descriptors for each metric.
@@
Avg 81.98 17.69 Avg 73.94 25.89
WavLM -Large +SE-ResNet Bright (明亮) 65.49 32.52 Bright (明亮) 49.33 48.27
@@
Avg 91.39 8.41 Avg 90.15 9.74
WavLM -Large +SE-ResNet Bright (明亮) 96.66 2.83 Bright (明亮) 85.61 12.96
@@ 5 Conclusion
This paper presents the CUHK-EE-DSP&STL systems submitted to the vTAD Challenge at NCMMSC 2025, designed to detect perceptual differences in voice timbre attributes through pairwise utterance comparison. Our approach integrates WavLM-Large embeddings with attentive statistical pooling for robust speaker representation, followed by two variants of Diff-Net, i.e., FFN and SE-ResFFN, for attribute intensity comparison. The systems achieved second place in the ‘Unseen’ track and fourth place in the ‘Seen’ track, demonstrating competitive performance across both generalisation and speaker-specific scenarios.

### v2

@@ Abstract
This paper presents the Voice Timbre Attribute Detection (vTAD) systems developed by the Digital Signal Processing & Speech Technology Laboratory (DSP&STL) of the Department of Electronic Engineering (EE) at The Chinese University of Hong Kong (CUHK) for the 20th National Conference on Human-Computer Speech Communication (NCMMSC 2025) vTAD Challenge. The proposed systems leverage WavLM-Large embeddings with attentive statistical pooling (ASTP) to extract robust speaker representations, followed by two variants of Diff-Net, i.e., Feed-Forward Neural Network (FFN) and Squeeze-and-Excitation-enhanced Residual FFN (SE-ResFFN), to compare timbre attribute intensities between utterance pairs. Experimental results demonstrate that the WavLM-Large+FFN system generalises better to unseen speakers, achieving 77.96% accuracy and 21.79% equal error rate (EER), while the WavLM-Large+SE-ResFFN model excels in the ‘Seen’ setting with 94.42% accuracy and 5.49% EER. These findings highlight a trade-off between model complexity and generalisation, and underscore the importance of architectural choices in fine-grained speaker modelling. Our analysis also reveals the impact of speaker identity, annotation subjectivity, and data imbalance on system performance, pointing to future directions for improving robustness and fairness in timbre attribute detection.
@@ 1 Introduction
The design of CUHK-EE systems entering the vTAD Challenge leverages a large-scale speech self-supervised learning (SSL) representation model, namely the WavLM-Large [3], to extract robust voice features from speech signals. The voice features are refined using attentive statistical pooling (ASTP) [4] before being passed into a comparison network, referred to as the Diff-Net. Two architectural variants of Diff-Net are investigated: standard Feed-Forward Neural Network (FFN) and Deep Residual FFN enhanced with Squeeze-and-Excitation blocks (SE-ResFFN) [5, 6, 7].
@@ 2.2 System Overview
The Diff-Net module outputs a vector of dimension $N$. Each element of the output vector represents a specific timbre attribute. A sigmoid function is applied to the output of the Diff-Net to produce the prediction vector $\hat{\rm y}$. The $n$-th element of $\hat{\rm y}$ (n=1,2,,34n=1,2,...,34) gives the likelihood that ${\mathcal{O}}_{\rm B}$ has a stronger intensity than ${\mathcal{O}}_{\rm A}$ in the $n$-th timbre descriptor.
@@ 2.4 Diff-Net
Figure 3 depicts the architectures of the two variants of Diff-Net investigated in our study. The first variant, illustrated as in Figure 3 (a), is an enhanced version of the traditional FFN [1, 2]. It consists of four sequential FFN blocks with hidden dimensions of $[512,256,128,64]$. Each block comprises a fully connected (FC) layer, followed by batch normalisation (BN), a ReLU activation function, and a dropout layer with a dropout rate of 0.3. After the final FFN block, an additional FC layer is applied to produce predictions for all 34 timbre attributes.
The second variant, SE-ResFFN, is inspired by the squeeze-and-excitation ResNet (SE-ResNet) architecture, which has shown strong performance in SV tasks [5, 7]. As illustrated in Figure 3 (b), (c), and (d), this model incorporates four SE-ResNet blocks with hidden dimensions of $[1024,1024,512,256]$. Following these blocks, a BN layer is applied, and the resulting features are passed through two FC layers with hidden dimensions $[192,64]$, interleaved with a ReLU activation. A final FC layer then outputs the predictions for the 34 timbre attributes.
@@ 3.1 Datasets
The VCTK-RVA dataset [12] is used for training the downstream Diff-Net modules (i.e., FFN and SE-ResFFN) and performance evaluation while the core component of the speaker encoder, the pre-trained WavLM-Large, is frozen in our training procedures. VCTK-RVA provides the annotation of 17 timbre descriptors for male and female speakers, resulting in a total of 34 distinct timbre attributes. The dataset was partitioned into a training set and two evaluation sets (‘Seen’ and ‘Unseen’), following the official challenge protocol. The training set includes 78 speakers (29 male and 49 female) forming 3,408 annotated speaker pairs and 136,320 utterance pairs. For evaluation, the ‘Seen’ set consists of 76 speakers who also appear in the training set but with different utterances, comprising 235 speaker pairs and 94,000 utterance pairs, while the ‘Unseen’ set contains 23 entirely new speakers forming 229 speaker pairs and 91,600 utterance pairs. The descriptors characterise different aspects of perceived voice quality, such as bright, thin, coarse, magnetic, shrill, and husky. The annotations are derived from human judgment, reflect subjective interpretations of vocal timbre. The subjectivity introduces variability in the data, which poses challenges for model consistency and generalisation to unseen speakers and utterances.
@@ 3.2 Training Settings
Model training was conducted over 10 epochs with a batch size of 16, using the AdamW optimiser with a learning rate of 1e41e-4 and a weight decay of 0.01. A cosine annealing scheduler was employed to adjust the learning rate dynamically throughout training. The loss function used was binary cross-entropy with sample-wise reduction, and model validation was performed after each epoch. To ensure reproducibility, all experiments were initialised with a fixed random seed (42).
@@
Table 1: The results on the ‘Unseen’ and ‘Seen’ test sets.
@@
WavLM-Large +SE-ResFFN 74.90 25.17
Seen WavLM-Large +FFN 90.77 9.08
WavLM-Large +SE-ResFFN 94.42 5.49
@@ 4 Results and Analysis
To further probe model robustness, we conducted small-scale experiments using speaker-disjoint training splits and observed considerable variation in results depending on the speaker composition. This, along with the inherent subjectivity in manual timbre labelling, suggests that model performance is sensitive to both speaker identity and annotation consistency. Additionally, the dataset [12] exhibits imbalance in descriptor frequency and gender representation–some attributes are under-represented, and female speakers dominate the data, yet male speakers often yield better results. These factors collectively highlight the need for more balanced data and refined annotation practices to improve generalisation and fairness. Detailed results for the ‘Unseen’ and ‘Seen’ test sets, broken down by gender and descriptor, are listed in Table 2.
Table 2: Evaluation results of our proposed systems on the ‘Unseen’ and ‘Seen’ test sets. The rows ‘Average’ are obtained by averaging the results across all the descriptors for each metric.
@@
Average 81.98 17.69 Average 73.94 25.89
WavLM -Large +SE-ResFFN Bright (明亮) 65.49 32.52 Bright (明亮) 49.33 48.27
@@
Average 91.39 8.41 Average 90.15 9.74
WavLM -Large +SE-ResFFN Bright (明亮) 96.66 2.83 Bright (明亮) 85.61 12.96
@@ 5 Conclusion
This paper presents the CUHK-EE systems submitted to the vTAD Challenge at NCMMSC 2025, designed to detect perceptual differences in voice timbre attributes through pairwise utterance comparison. Our approach integrates WavLM-Large embeddings with attentive statistical pooling for robust speaker representation, followed by two variants of Diff-Net, i.e., FFN and SE-ResFFN, for attribute intensity comparison. The systems achieved second place in the ‘Unseen’ track and fourth place in the ‘Seen’ track, demonstrating competitive performance across both generalisation and speaker-specific scenarios.